OpenStudy (anonymous):

A two-day conference is going to be held in a city. The leading newspaper in the city says there is a 10% chance of rain on the first day of the conference and a 40% chance of rain on the second day. If the newspaper’s figures are correct, the chance that it rains in the city during the conference is: at least 0%, but we can’t get a better lower bound with the information given. at least 10% but it might be less than 40%. at least 40%.

OpenStudy (anonymous):

If the newspaper’s figures are correct, the chance that it rains in the city during the conference is at most 10%. at most 40%. at most 50%. at most 100%; that’s the best upper bound we can get based on the information given.

OpenStudy (kropot72):

Assuming the events 'rains on first day' and 'rains on second day' are independent, the probability of not raining on either day is given by \[P(no\ rain\ either\ day)=(1-0.1)\times (1-0.4)=0.9\times 0.6=0.54\] Therefore \[P(rains\ during\ conference)=1-0.54=0.46\]

OpenStudy (anonymous):

can any one post the corect ans for the above question

OpenStudy (anonymous):

A two-day conference is going to be held in a city. The leading newspaper in the city says there is a 10% chance of rain on the first day of the conference and a 40% chance of rain on the second day. PROBLEM 6A This Problem is worth 1 Point If the newspaper’s figures are correct, the chance that it rains in the city during the conference is: at least 0%, but we can’t get a better lower bound with the information given. at least 10% but it might be less than 40%. at least 40%.

OpenStudy (anonymous):

Hi, 6a) at least 40% it's right answer good luck!

OpenStudy (anonymous):

I am with a group of five of my friends. A box contains 8 pieces each of milk chocolate, white chocolate, and dark chocolate. The box is passed around the six of us, with each person taking 4 pieces. Assume that each person chooses at random without replacement from the available pieces. I am the last person to whom the box is passed. Find the chance that I pick 4 dark chocolates.

OpenStudy (anonymous):

plsssss ans